# Incorporating Judgment Prediction into Legal Case Retrieval via Law-aware Generative Retrieval

### v1

@@ Abstract.
Legal case retrieval and judgment prediction are crucial components in intelligent legal systems. In practice, determining whether two cases share the same charges through legal judgment prediction is essential for establishing their relevance in case retrieval. However, current studies on legal case retrieval merely focus on the semantic similarity between paired cases, ignoring their charge-level consistency. This separation leads to a lack of context and potential inaccuracies in the case retrieval that can undermine trust in the system’s decision-making process. Given the guidance role of laws to both tasks and inspired by the success of generative retrieval, in this work, we propose to incorporate judgment prediction into legal case retrieval, achieving a novel law-aware Generative legal case retrieval method called Gear. Specifically, Gear first extracts rationales (key circumstances and key elements) for legal cases according to the definition of charges in laws, ensuring a shared and informative representation for both tasks. Then in accordance with the inherent hierarchy of laws, we construct a law structure constraint tree and assign law-aware semantic identifier(s) to each case based on this tree. These designs enable a unified traversal from the root, through intermediate charge nodes, to case-specific leaf nodes, which respectively correspond to two tasks. Additionally, in the training, we also introduce a revision loss that jointly minimizes the discrepancy between the identifiers of predicted and labeled charges as well as retrieved cases, improving the accuracy and consistency for both tasks. Extensive experiments on two datasets demonstrate that Gear consistently outperforms state-of-the-art methods in legal case retrieval while maintaining competitive judgment prediction performance.
Legal case retrieval, Generative Retrieval, Legal Judgment Prediction
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
Legal case retrieval, which seeks to retrieve relevant precedent cases based on a usually undecided case, stands as a pivotal component in intelligent legal systems. This task relies on specialized legal knowledge encoded in laws, with the goal of providing essential evidence to legal practitioners and supporting their decision-making processes.
Existing studies on legal case retrieval (Ma et al., 2021; Xiao et al., 2021a; Chalkidis et al., 2020) have been focused on the semantic similarity between paired cases, for example,  Li et al. (2023) utilize the structural information contained in legal cases to conduct fine-grained semantic interactions between the query and candidate cases. Although they have achieved impressive performances, they ignore the fact that legal case retrieval and judgment prediction [^1] Legal judgment prediction aims to predict the relevant law article, charge, and term of penalty for a given query case. In this work, we focus on the charge prediction. Other sub-tasks are easily incorporated into our proposed method. are closely intertwined (Shao et al., 2023; Zhang and Dou, 2023). In particular, the charge-level consistency between the query and retrieved case, a direct product of the judgment prediction, is crucial for relevance estimation in legal case retrieval. As shown in Figure 1, it provides valuable contextual information and effective relevance signals. Moreover, incorporating judgment prediction into case retrieval improves the transparency (Olteanu et al., 2021; Kim et al., 2020; Baan et al., 2019) of the decision-making process. This integration aids legal practitioners in tracing the retrieved cases from the perspective of charges, thereby making the legal decisions trustworthy.
Figure 1. A pair of relevant cases from the LeCaRDv2 dataset. The applicable charges (underlined) and their corresponding rationales (red and blue) provide effective relevance signals.
Despite the aforementioned benefits, there remain the following challenges in incorporating judgment prediction into case retrieval. Firstly, legal case retrieval and judgment prediction are usually formulated as two distinct machine learning problems–ranking and classification. It is difficult for one ranking model to predict the applicable charges for legal cases and leverage the charge prediction to enhance ranking. Secondly, legal cases are inherently lengthy and complicated. As shown in Figure 1, only the rationales (key circumstances and key elements)  (Ma et al., 2021; Yu et al., 2022) of the case provide effective signals for relevance estimation. It is expected to represent each case as a shared and informative representation for both tasks, but existing methods often encode legal cases in a coarse-grained granularity way (Chalkidis et al., 2020; Xiao et al., 2021a). Thirdly, both tasks rely on specialized law knowledge (Sun et al., 2022; Liu et al., 2023c), an appropriate way that effectively incorporates law expertise to guide the prediction of both tasks remains a concern, given that current researches (Chalkidis et al., 2020; Xiao et al., 2021a; Zhang et al., 2023a; Li et al., 2023) solely focus on text-level semantics for predictions.
Facing the above challenges and inspired by the success of generative retrieval (Tay et al., 2022; Wang et al., 2022), we propose a novel generative legal case retrieval method, namely Gear, to incorporate judgment prediction into legal case retrieval. As illustrated in Figure 2, Gear unifies these two tasks into a tree-based generative retrieval framework. The insight of Gear is to map the query case to a set of law-aware semantic identifiers where each identifier represents a precedent case relevant to the query and also reflects applicable charges. Specifically, we first construct a corpus based on the applicable laws [^2] We consider the Criminal Law of the People’s Republic of China in this work. and subsequently extract rationales at word- and sentence-level for each case according to this corpus. It effectively filters out the noise and ensures the case representations are shared and informative to both tasks. Then, we create the law structure constraint tree based on the inherent hierarchy of laws (Chapter-Section-Article) , considering that both tasks are learned with the guidance of laws. Thus, we can consider situations where each case may involve single or multiple charges and assign the law-aware semantic identifier(s) to each case given this tree, making the identifier reflect its charges. In this way, the retrieval is equivalent to a traversal from the root, through intermediate charge nodes, to case-specific leaf nodes, which respectively correspond to two tasks. To improve the accuracy and consistency for both tasks, we develop a revision loss that jointly minimizes the discrepancy between predicted and labeled charges/ retrieved cases. Furthermore, to assess the performance of judgment prediction in case retrieval, we introduce a novel metric called coverage. This metric gauges the charge-level consistency between the query and retrieved cases.
@@
(1) To the best of our knowledge, this is the first work that incorporates judgment prediction into legal case retrieval, addressing the challenge of charge-level consistency between the query case and retrieved cases in legal search. Our work also improves the transparency of the legal decision-making process.
(2) We propose a novel model, namely Gear, to unify legal case retrieval and judgment prediction under the generative retrieval framework. We explicitly leverage the law knowledge to extract rationales from cases, assign cases the law-aware hierarchical identifies, and formulate the prediction as a traversal on the law structure constraint tree. We also propose a revision loss to jointly improve the accuracy and consistency of both tasks.
(3) We conduct extensive experiments on two public datasets and propose an evaluation metric for assessing the charge-level consistency between the query and retrieved cases. Experimental results indicate Gear not only achieves state-of-the-art performance in legal case retrieval but also maintains competitive judgment prediction performance.
@@ 2.1. Task Formulation
In this work, we target on the task of legal case retrieval. Suppose that we have a set of collected samples $\mathcal{D}=\{(q,\mathcal{C},\mathcal{R})\}$. For each data instance, $q$ represents a undecided query case submitted by the legal practitioner; $\mathcal{C}=\{c_{1},c_{2},\cdots,c_{N}\}$ represents candidate cases (precedents) pool with size $N\in\mathbb{N}^{+}$; $\mathcal{R}$ represents the labeled relevant case set from $\mathcal{C}$ given the query case $q$. Unlike previous studies that only predict to retrieve $\mathcal{R}$ from $\mathcal{C}$ given $q$, in this work, we instead unify judgment prediction and case retrieval into a generative retrieval framework, and thus aims at learning a ranking function $f:q\times\mathcal{C}\rightarrow\mathcal{R}\times\mathcal{E}$, where $\mathcal{E}$ denotes the set of applicable charges corresponding to $q$.
2.2. The Principle of Our Method
To retrieve relevant cases and predict their applicable charges, we develop a novel learning method from the viewpoint of generative retrieval (GR). Essentially, GR is an emerging new retrieval paradigm, that adopts a deep neural network to perform the end-to-end retrieval in contrast to traditional the index-then-retrieve paradigm. GR models consist of two major steps to directly generate identifier strings of documents as the retrieval target. In the first indexing step that focuses on memorizing the information about each document, GR models take each original document $c$ as input and generate its identifier ${id}^{c}$ as output in a straightforward Seq2Seq fashion. The model is trained with the standard language model objective with the teacher forcing:
(1) $\mathcal{L}_{i}=\sum_{c\in\mathcal{C}}\log P({id}^{c}|c),$
where ${id}^{c}$ is represented by the atomic identifier or the semantic identifier. Atomic identifiers are generated by assigning each document an arbitrary integer, while semantic identifiers are generated according to the hierarchical clustering algorithm employed over all the documents. In the second retrieval phase, GR models associate each $q$ to its relevant document $r$ through an auto-regressive generation:
@@ 2.2. The Principle of Our Method
where ${id}^{r}$ denotes the identifier of $r$. ${id}^{r}$ is also represented by the atomic identifier or the semantic identifier. As such, once a GR model is learned, it can be used to retrieve candidate documents for a test query in an end-to-end manner using beam search.
As aforementioned in Section 1, applying GR to our legal scenario faces several challenges including the lengthy and complicated content, the law-aware semantic identifier, and the training objective for enhancing both tasks. Therefore, when incorporating legal judgment prediction into case retrieval in the above GR framework, the critical learning tasks become: (1) extract rationales instead of using original documents to form the input for GR (Section 2.3); (2) create informative law-aware identifiers for each case based on the hierarchical structure of laws (Section 2.4); (3) develop a training objective to explicitly ensure charge-level and case-level consistency between predictions and labels(Section 2.5).
Figure 3. The proposed three modules of Gear. $f_{R}$ extract rationales from legal cases; $f_{T}$ assigns hierarchical identifiers to each case and constrain the decoding; $\mathcal{L}_{c}$ jointly optimizes case retrieval and judgement prediction.
@@ 2.3. Rationale Extraction
To filter the noise and provide shared and informative representations for legal cases, based on the set of laws $L$, we devise a module called $f_{R}$ to extract rationales $E$ instead of using the original document $doc\in\{q,c\}$ as the input of Gear:
@@
As shown in Figure 3, to achieve this goal, we first leverage the guidance of laws to construct a corpus $B$, collecting law-based keywords and embeddings. Specifically, the corpus consists of three parts $B=\{B_{1},B_{2},B_{3}\}$. $B_{1}$ is constructed according to the lexical variants of all charges names in $L$, for example, in terms of the charge “crime of forges the seals of a company, enterprise, institution or a people’s organization”(translated from Chinese), we split the charge name and put all keywords and their lexical variants to the corpus after the stop words are removed. $B_{2}$ is constructed similarly to the $B_{1}$. We put all keywords in the definition of laws and the feedback from prompting ChatGPT [^3] https://openai.com/chatgpt, the prompts are provided in Appendix B.2. to the corpus after the stop words are removed. $B_{3}$ is constructed by collecting BERT (Devlin et al., 2018; Chalkidis et al., 2020) embeddings [^4] https://github.com/thunlp/OpenCLaP. for the definition of each law.
Once the corpus is collected, we compute the multiple-level scores and extract $E$ for each case based on three corpora as follows. At the word level, we split the case and respectively select the top-$k_{1}$ (top-$k_{2}$) keywords $E_{w_{1}}$ ($E_{w_{2}}$) from $B_{1}$ ($B_{2}$) as follows:
@@
where $w$ is the word from $s$. Please note that since query cases are typically undecided with no charges, we extract rationales using the method described above. On the other hand, candidate cases are precedents with decided charges, thus, when extracting the rationales for them, we shrunk $B$ into a corpus constructed based on its corresponding charges.

### v2

# Explicitly Integrating Judgment Prediction with Legal Document Retrieval: A Law-Guided Generative Approach

@@ Abstract.
Legal document retrieval and judgment prediction are crucial tasks in intelligent legal systems. In practice, determining whether two documents share the same judgments is essential for establishing their relevance in legal retrieval. However, existing legal retrieval studies either ignore the vital role of judgment prediction or rely on implicit training objectives, expecting a proper alignment of legal documents in vector space based on their judgments. Neither approach provides explicit evidence of judgment consistency for relevance modeling, leading to inaccuracies and a lack of transparency in retrieval. To address this issue, we propose a law-guided method, namely GEAR, within the generative retrieval framework. GEAR explicitly integrates judgment prediction with legal document retrieval in a sequence-to-sequence manner. Specifically, given the intricate nature of legal documents, we first extract rationales from documents based on the definition of charges in law. We then employ these rationales as queries, ensuring efficiency and producing a shared, informative document representation for both tasks. Second, in accordance with the inherent hierarchy of law, we construct a law structure constraint tree and represent each candidate document as a hierarchical semantic ID based on this tree. This empowers GEAR to perform dual predictions for judgment and relevant documents in a single inference, i.e., traversing the tree from the root through intermediate judgment nodes, to document-specific leaf nodes. Third, we devise the revision loss that jointly minimizes the discrepancy between the IDs of predicted and labeled judgments, as well as retrieved documents, thus improving accuracy and consistency for both tasks. Extensive experiments on two Chinese legal case retrieval datasets show the superiority of GEAR over state-of-the-art methods while maintaining competitive judgment prediction performance. Moreover, we validate the effectiveness of GEAR on a French statutory article retrieval dataset, reaffirming its robustness across languages and domains.
Legal Document Retrieval, Generative Retrieval, Legal Judgment Prediction
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 14–18, 2024; Washington, DC, USAbooktitle: Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’24), July 14–18, 2024, Washington, DC, USAdoi: 10.1145/3626772.3657717isbn: 979-8-4007-0431-4/24/07ccs: Information systems Information retrievalccs: Applied computing Law
@@ 1. Introduction
Legal document retrieval and judgment prediction are fundamental components in intelligent legal systems. The former entails the retrieval of relevant legal documents (cases or statutory articles) give a query. On the other hand, the latter seeks to predict the outcomes or judgments rendered in legal cases, such as applicable charges, term-of-penalties, etc.
@@
These two tasks are closely intertwined (Shao et al., 2023a, b; Zhang and Dou, 2023) in practice. From the retrieval side, determining whether two documents share the same judgments is essential for establishing their relevance. Regrettably, most of the existing studies (Ma et al., 2021; Xiao et al., 2021a; Chalkidis et al., 2020; Sun et al., 2024; Yu et al., 2022a) of legal document retrieval frequently overlook the significance of judgment prediction and merely focus on the text-level semantic similarity. Recently, Li et al. (2023a) introduced an implicit training objective that uses the fact description of the legal document to predict its judgment, expecting a proper alignment of legal documents in vector space based on their judgments. While these studies show effectiveness in retrieval performance, they fail to provide explicit evidence of judgment consistency for relevance modeling. Consequently, this limitation leads to inaccuracies and a lack of transparency (Olteanu et al., 2021; Kim et al., 2020; Baan et al., 2019). It is because their legal relevance reasoning especially regarding judgment remains unclear, and we cannot trace back the decision-making process based on the retrieval results.
Therefore, we aim to explicitly integrate judgment prediction with legal document retrieval. However, there remain the following challenges to achieve our goal. Firstly, legal document retrieval and judgment prediction are usually formulated as two distinct machine learning problems–retrieval and classification. It is difficult for one retrieval model to predict the applicable judgment for legal cases and in turn leverage the judgment prediction to enhance retrieval. Secondly, legal documents are inherently lengthy and complicated. It results in the retrieval efficiency issue and hiders to represent each document as a shared and informative representation for both tasks. Thirdly, both tasks rely on specialized law knowledge (Sun et al., 2022; Liu et al., 2023b), an appropriate way that effectively injects law expertise to guide the prediction of both tasks remains a concern.
Facing the above challenges, we propose a novel law-guided generative legal document retrieval method, namely GEAR. GEAR explicitly integrates judgment prediction with legal document retrieval in a sequence-to-sequence (Seq2Seq) manner as illustrated in Figure 1. The insight of GEAR lies in formulating the retrieval process into the generation of law-aware semantic IDs, where each ID not only represents a document relevant to the query but also reflects its applicable judgments. Specifically, we first construct a corpus based on the definition of charges in law [^1] In this paper, we take laws in China and Belgium as examples. It is worth noting that GEAR can readily accommodate laws from various other countries. and subsequently extract rationales (Yu et al., 2022b; Sun et al., 2023) representing the key elements/ circumstances (Ma et al., 2021; Shao et al., 2023b) at the word- and sentence-level for each document according to this corpus. We employ these rationales instead of raw documents as queries for generation. This strategy not only serves to effectively filter out the noise of legal documents, ensuring generation efficiency but also renders the rationales shared and informative for both tasks. Then, we create the law structure constraint tree based on the inherent hierarchy of law (e.g. Chapter-Section-Article), considering that both tasks are learned with the guidance of law. Given this tree, we assign legal documents hierarchical semantic IDs, with the IDs reflecting their judgments, for example, the ID “0-2-5-269-809” indicates the document named 809 falls under Article 269 of Chapter 5 of Section 2. In this way, the generation of these IDs is equivalent to traversing the tree from the root through intermediate judgment nodes, to document-specific leaf nodes. It makes GEAR capable of showing the legal reasoning process and performing dual predictions for judgment and relevant documents in a single inference. To further improve the accuracy and consistency of both tasks, we devise a novel training objective called the revision loss. This loss aligns with the hierarchy of the tree and jointly minimizes the discrepancy between predicted and labeled judgments/ retrieved cases. Extensive experiments on two Chinese legal case retrieval datasets show the consistent superiority of GEAR over state-of-the-art methods while maintaining competitive judgment prediction performance. We also validate the effectiveness of GEAR on a French statutory article retrieval dataset, reaffirming its generalization ability.
@@
(1) To the best of our knowledge, this is the first work that explicitly integrates judgment prediction with legal document retrieval. Our method is capable of showing the legal reasoning process and performing dual predictions for both tasks in a single inference. It improves the transparency of the legal decision-making.
(2) We propose a novel law-guided generative model, namely GEAR. We explicitly leverage the law knowledge to extract rationales from legal documents, assign them the law-aware hierarchical IDs, and formulate the prediction as a traversal on the law structure constraint tree. We also propose the revision loss to jointly improve the accuracy and consistency of both tasks.
(3) We conduct extensive experiments on three public datasets of legal document retrieval in two languages. The results indicate GEAR not only achieves state-of-the-art performance in legal case and statutory article retrieval but also maintains competitive judgment prediction performance.
2. Related Work
2.1. Legal Document Retrieval
Legal document retrieval is a long-standing research topic in the field of information retrieval. In the early exploration, researchers (Moens, 2001; Zeng et al., 2005; Klein et al., 2006; Saravanan et al., 2009; Bench-Capon et al., 2012) made efforts to inject the legal knowledge to the retrieval through the decomposition of legal issues and involving the ontology. In recent years, deep learning demonstrated its effectiveness in exploring document semantics. One representative line of works focused on the network-based precedents methods tailored for the common law systems. For example, Minocha et al. (2015) leveraged the Precedent Citation Network (PCNet) to predict the relevance based on whether the sets of precedent citations occur in the same cluster. Bhattacharya et al. (2020) proposed Hier-SPCNet to capture all domain information inherent in laws and precedents. The other line of works judged the relevance between the query and candidate document according to their text-level similarity. One such method was BERT-PLI (Shao et al., 2020). It divided the legal document into several paragraphs and used BERT (Devlin et al., 2018) to obtain the similarity between the paragraphs. Lawformer (Xiao et al., 2021b) was another text-based method. It used millions of Chinese criminal and civil case documents to pre-train a Longformer (Beltagy et al., 2020) model. Despite impressive performance, these works overlook the significance of judgment prediction and merely focus on the text-level similarity, resulting in sub-optimal and unreliable results.
Recently, Li et al. (2023a) introduced an implicit training objective that uses the fact description of the legal document to predict its judgment, expecting a proper alignment of legal documents in vector space based on their judgments. It fails to provide explicit evidence of judgment consistency for relevance modeling, leading to inaccuracies and a lack of transparency.
2.2. Generative Retrieval
Generative retrieval has recently emerged as a promising direction for document retrieval. These methods assign semantic IDs to documents and utilize language models for ID generation. It enables an end-to-end retrieval in contrast to the traditional index-then-retrieve paradigm. Various methods have been introduced to generate semantic document IDs. For example,  Cao et al. (2021) introduced a Seq2Seq system to conduct entity retrieval. They first represented documents as unique names that are composed of entity names. Then they used a auto-regressive generation model to generate the unique names of these entities based on contextual information.  Tay et al. (2022) proposed the differentiable search index (DSI) paradigm, which is an auto-regressive generation model to perform ad-hoc retrieval tasks. The input of the model was a natural language query and the model regressively generated documents’ ID strings that are relevant to the given query.  Wang et al. (2022) proposed a novel method NCI, which used a tailored prefix-aware weight-adaptive decoder to optimize the retrieval performance. Ultron (Zhou et al., 2022) leveraged document titles and substrings as IDs to enrich the semantic information of IDs. To mitigate data distribution mismatch that occurs between the indexing and the retrieval phases,  Zhuang et al. (2022) proposed DSI-QG, which adopted a query generation model with a cross-encoder to generate and select a set of relevant queries.
Existing studies focused on the general domain, lacking specific designs for legal documents and the integration of law knowledge.
@@ 3.1. Task Formulation
In this work, we target on legal document retrieval including legal case retrieval (LCR) and statutory article retrieval (SAR). Suppose that we have a set of collected samples $\mathcal{D}=\{(q,\mathcal{C},\mathcal{R})\}$. For each data instance, $q$ is the query representing an undecided legal case submitted by the legal practitioner in LCR, a legal question in SAR; $\mathcal{C}=\{c_{1},c_{2},\cdots,c_{N}\}$ with size $N\in\mathbb{N}^{+}$ is the candidate precedent case set in LCR, the statutory article pool in SAR; $\mathcal{R}$ represents the labeled relevant case/ statutory article set from $\mathcal{C}$ given the query $q$. Unlike previous studies that only predict to retrieve $\mathcal{R}$ from $\mathcal{C}$ given $q$, in this work, we instead unify judgment prediction and document retrieval into a generative retrieval framework, and thus aim at learning a retrieval function $f:q\times\mathcal{C}\rightarrow\mathcal{R}\times\mathcal{E}$, where $\mathcal{E}$ denotes the set of applicable judgment corresponding to $q$.
3.2. Overall Framework
To learn $f$ and explicitly integrate judgment prediction with legal document retrieval, we develop GEAR, a novel law-guided generative approach from the viewpoint of generative retrieval. Essentially, given a query document, GEAR adopts a language model to perform the Seq2Seq generation where the retrieved documents are represented as semantic IDs. Following the practice of (Tay et al., 2022; Wang et al., 2022; Zhuang et al., 2022; Zhou et al., 2022), GEAR consists of two major steps to directly generate IDs of documents as the retrieval target. In the first indexing step that focuses on memorizing the information about each document, our GEAR takes each document $c$ as input and generates its ID ${id}^{c}$ as output. The model is trained with the standard language model objective with the teacher forcing:
(1) $\mathcal{L}_{i}=\sum_{c\in\mathcal{C}}\log P({id}^{c}|c).$
In the second retrieval phase, GEAR models associate each $q$ to its relevant document $r$ through an auto-regressive generation:
@@ 3.2. Overall Framework
where ${id}^{r}$ denotes the ID of $r$. As such, once a GEAR model is trained, it can be used to retrieve candidate documents for a test query in an end-to-end manner using beam search.
As aforementioned in Section 1, there are several challenges in the legal domain. For explicitly integrating judgment prediction with legal document retrieval within GEAR, the critical learning tasks become: (1) extract rationales instead of using raw documents to form the input for generation (Section 3.3); (2) create informative law-aware IDs for each document based on the hierarchical structure of law (Section 3.4); (3) develop a training objective to explicitly ensure judgment-level and document-level consistency between predictions and labels (Section 3.5).
Figure 2. The proposed three modules of GEAR. $f_{R}$ extract rationales from legal documents; $f_{T}$ assigns hierarchical IDs to each document and constrain the decoding; $\mathcal{L}_{c}$ jointly optimizes document retrieval and judgement prediction.
@@ 3.3. Rationale Extraction
To ensure efficiency and provide shared and informative representations for legal documents, based on the set of laws $L$, we devise a module called $f_{R}$ to extract rationales $E$ instead of using the raw document $doc\in\{q,c\}$ as the input of GEAR:
@@
As illustrated in Figure 2, we first leverage the guidance of law to construct a corpus $B=\{B_{1},B_{2},B_{3}\}$, collecting law-based keywords $B_{1},B_{2}$ and embeddings $B_{3}$. The keyword set $B_{1}$ is constructed according to the lexical variants of all charge names in $L$. For example, in terms of the charge “crime of forges the seals of a company, enterprise, institution or a people’s organization”(translated from Chinese), we split the charge name and remove the stop words, then add the rest and their lexical variants to $B_{1}$. The keyword set $B_{2}$ is constructed similarly to the $B_{1}$. We split the definitions of charges in law and remove the stop words then add the rest and their lexical variants to $B_{2}$. For the augmentation purpose, we employ the definitions of charges in law as prompts for the large language model [^2] We use ChatLaw (Cui et al., 2023) for Chinese data. Since we have not found a suitable LLM for the Belgian legal domain, we omit this step for this data. (LLM) designed for the legal domain. After obtaining feedback from LLM, we remove stop words and incorporate the results into $B_{2}$. To avoid the hallucination issue, we engage legal experts to manually assess the quality of augmentations to ensure the effectiveness of $B_{2}$. $B_{3}$ is constructed by collecting Legal-BERT (Devlin et al., 2018; Chalkidis et al., 2020) embeddings for the definition of each charge in law.
Once the corpus is collected, we compute the multiple-level scores and extract $E$ for each document based on three corpora as follows. At the word level, we split the document and respectively select the top-$k_{1}$ (top-$k_{2}$) keywords $E_{w_{1}}$ ($E_{w_{2}}$) from $B_{1}$ ($B_{2}$) as follows:
@@
where $w$ is the word from $s$. Please note that since query documents are typically undecided i.e., without labeled applicable charges, we extract rationales using the method described above. For the candidate precedent documents, whose applicable charges are given, we shrunk $B$ into a corpus constructed based on their corresponding labeled charges.